OpenStudy (anonymous):

3m^2-2m=2, its a lovely equation that must be solved via quadratic formula, help is much appreciated :3 =D

OpenStudy (anonymous):

the stupid math program erased the other one :( so i got a new one, this one

OpenStudy (anonymous):

\[3m^2 - 2m - 2 = 0\] a = 3 b = -2 c = -2

OpenStudy (anonymous):

three answers?? o.o

OpenStudy (anonymous):

\[\frac{2 \pm \sqrt{(-2)^2 - 4(3)(-2)}}{6}\]

OpenStudy (anonymous):

Nono.. thats the numbers you plug into the quadratic formula

OpenStudy (anonymous):

\[\frac{2 \pm \sqrt{28}}{6}\]\[\frac{2 \pm 2 \sqrt{7}}{6}\] \[\frac{ 1 \pm \sqrt{7}}{3}\]

OpenStudy (anonymous):

woops lol thanks ^^; i was so tired

OpenStudy (anonymous):

You're Welcome! :)

OpenStudy (anonymous):

:D